OpenStudy (anonymous):

Evaluate the difference quotient f(x) = x + 3 x + 1 , f(x) − f(1) x − 1

OpenStudy (anonymous):

maybe \[\frac{f(x)-f(1)}{x-1}\]?

OpenStudy (anonymous):

i dont know ive been looking for two days now and im still clueless

OpenStudy (anonymous):

i mean is that what you have to compute? it is not exactly clear from what you wrote if that is the what you need, we can do it no problem

OpenStudy (anonymous):

dang just realized i had the wrong thing wrote F(x) = (x+3)/(x+1) , (f(x) - f(10)) / (x-1)

OpenStudy (anonymous):

maybe \[\frac{f(x)-f(1)}{x-1}\]?

OpenStudy (anonymous):

good lord i give up at typing, its to late F(x) = (x+3)/(x+1) , (f(x) - f(1)) / (x-1)

OpenStudy (anonymous):

ok ready?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

first we need \(f(1)\) which is easy enough, since \[f(x)=\frac{x+3}{x+1}\] we know \[f(1)=\frac{1+3}{1+1}=2\]

OpenStudy (anonymous):

so actual job is to compute \[\frac{\frac{x+3}{x+1}-2}{x-1}\]

OpenStudy (anonymous):

now there is no avoiding the algebra, we have to actually do the subtraction in the numerator, unless you want to multiply top and bottom by \(x+1\) to clear the fraction lets go ahead and subtract, working only in the numerator

OpenStudy (anonymous):

\[\frac{x+3}{x+1}-2=\frac{x+3-2(x+1)}{x+1}\] \[=\frac{x+3-2x-2}{x+1}=\frac{-x+1}{x+1}\]

OpenStudy (anonymous):

now we get to \[\frac{\frac{-x+1}{x+1}}{x-1}=\frac{-x+1}{(x+1)(x-1)}\] and the usual miracle occurs, one factor in the denominator will cancel since \[\frac{-x+1}{x-1}=-1\] your final answer is \[\frac{-1}{x-1}\]

OpenStudy (anonymous):

my most sincere thankyou's i belive that is correct

OpenStudy (anonymous):

or if you prefer \[-\frac{1}{x+1}\]

OpenStudy (anonymous):

all steps are there, if any are confusing let me know

OpenStudy (anonymous):

yw